# Score Distillation via Reparametrized DDIM

### v1

@@ 1 Introduction
Based on our analysis, we propose an alternative score distillation algorithm dubbed Score Distillation via Inversion (SDI), closing the gap to DDIM. We obtain the conditional noise required for consistency of the denoising trajectories by inverting DDIM on each step of score distillation (fig. 5). This modification yields 3D objects with high-quality textures consistent with the 2D diffusion model. Moreover, in 2D, our method closely approximates DDIM while preserving the incremental generation schedule of SDS (fig. 1d).
@@
  • We propose a new method titled Score Distillation via Inversion (SDI), which replaces the problematic random noise sampling in SDS with DDIM inversion and significantly improves 3D generation, closing the quality gap to samples from the 2D model.
@@ 2 Related work
Other papers propose heuristics that improve SDS. For example, [katzir2023noise, yu2023text] decompose the guidance terms and speculate about their relative importance. Empirically, visual quality can be improved by suppressing the denoising term with negative prompts [katzir2023noise] or highlighting the classification term [yu2023text]. [lin2023magic3d, tang2023make, wang2024prolificdreamer, chen2024it3d] use multi-stage optimization: they first train a volumetric representation and then extract a mesh or voxel grid to fine-tune geometry and texture. HiFA [zhu2023hifa] uses a combination of ad-hoc techniques: additional supervision in latent space, time annealing, kernel smoothing, $z$-variance regularization, and a pretrained mono-depth estimation network to improve the quality of single-stage NeRF generation. Rather than augmenting the SDS pipeline or relying on heuristics, we derive Score Distillation through the denoising process of DDIM and propose a simple modification of SDS that significantly improves 3D generation.
@@ 3 Background
Denote the parameters of a differentiable 3D shape representation by $\psi\in\mathbb{R}^{d}$ and differentiable rendering by a function $g(\psi,c):\mathbb{R}^{d}\times C\rightarrow\mathbb{R}^{N\times N}$ that returns an image given camera parameters $c\in C$. Intuitively, in each iteration, SDS samples $c$, renders the corresponding (image) view $g(\psi,c)$, perturbs it with $\epsilon\sim\mathcal{N}(0,I)$ to level $t\sim[0,1]$, and denoises it with $\epsilon_{\theta}^{t}$; the difference between the true and predicted noise is propagated to the parameters of the 3D shape. More formally, after sampling the camera view $c$ and randomly drawing a time $t$, SDS renders the volume and adds Gaussian noise $\epsilon$ to obtain a noisy image $x(t)=\sqrt{\alpha(t)}g(\psi,c)+\sqrt{1-\alpha(t)}\epsilon.$ Then, SDS updates the volume by using a gradient(-like) direction to update its parameters $\psi$:
@@ 4 Linking SDS to DDIM
In words, $x_{0}(t)$ is the noisy image at time $t$ denoised with a single step of noise prediction. Empirically, the evolution of $x_{0}(t)$ is similar to the evolution of $g(\psi,c)$—from blurry to sharp. In the left side of 4 we compare these processes. This similarity motivates us to rewrite eq. 1 in terms of $x_{0}(t)$, and to understand SDS as applying similar updates to the renderings of its 3D representation.
@@
Define $\kappa^{t}\bigl{(}x_{0}(t)\bigr{)}=\epsilon$ as a solution of this equation given $x_{0}(t)$. Then, we can write:
$\epsilon_{\theta}^{t}\bigl{(}x(t),y\bigr{)}=\kappa^{t}\bigl{(}x_{0}(t)\bigr{)}\textrm{\quad and\quad}x(t-\tau)=\sqrt{\alpha(t-\tau)}x_{0}(t)+\sqrt{1-\alpha(t-\tau)}\kappa^{t}\bigl{(}x_{0}(t)\bigr{)}\text{.}$ (9)
@@
$x_{0}(t\!-\!\tau)=x_{0}(t)\!-\!\sigma(t\!-\!\tau)\bigl{[}\underbrace{\epsilon_{\theta}^{t-\tau}\bigl{(}\overbrace{\sqrt{\alpha(t\!-\!\tau)}x_{0}(t)\!+\!\sqrt{1\!-\!\alpha(t\!-\!\tau)}\kappa^{t}\bigl{(}x_{0}(t)\bigr{)}}^{\text{$x_{0}$ noised with $\kappa^{t}$ to time $t-\tau$}},y\bigr{)}}_{\text{predicted noise}}\!-\!\underbrace{\kappa^{t}\bigl{(}x_{0}(t)\bigr{)}}_{\text{noise sample $\kappa^{t}$}}\!\!\!\bigr{]}\!\text{.}$ (10)
We can already see that the structure of eq. 10 is very similar to the SDS update rule in eq. 4. Note that the update direction in eq. 10 is the same as in the SDS update rule in eq. 4, where $\kappa^{t}$ plays the role of the random noise sample $\epsilon$. We could use it as a guidance for the 3D generative process in SDS by replacing $\epsilon$ in eq. 4 with $\kappa^{t}(x_{0}(t)).$ In practice, however, it is hard to solve eq. 8, as $\epsilon_{\theta}^{t}$ is high-dimensional and nonlinear. In an unconstrained 2D generation, $\kappa^{t}$ can be cached from a previous denoising step, matching the update step to DDIM exactly fig. 1(c). In 3D, however, this is impossible due to the simultaneous optimization of multiple views and projections to the space of viable 3D shapes. Below we show that a naïve approximation replacing $\kappa^{t}$ with a Gaussian yields SDS, and we will propose alternatives that are more faithful to the derivation above.
SDS as a special case. From eq. 10, to get a cleaner image, we need to bring the current image to time $t$ with noise sample $\kappa^{t}$, denoise the obtained image, and then subtract the difference between added and predicted noise from the initial image. A coarse approximation of $\kappa$ uses i.i.d. random noise $\kappa^{t}_{\textrm{SDS}}(x_{0}(t))\sim\mathcal{N}(0,I)$, matching the forward process by which diffusion adds noise. This choice of $\kappa^{t}_{\textrm{SDS}}$ precisely matches the update rule eq. 10 to the SDS guidance in eq. 4.
@@ 5 Score Distillation via Inversion (SDI)
As we have shown, SDS follows the velocity field of reparametrized DDIM in eq. 10, when $\kappa_{\textrm{SDS}}^{t}(x_{0}(t))$ is randomly sampled in each step. Our derivation, however, suggests that $\kappa^{t}_{\textrm{SDS}}(x_{0}(t))$ could be improved by bringing it closer to a solution of the fixed-point equation in eq. 8. Indeed, randomly sampling $\kappa^{t}_{\textrm{SDS}}$ as in Dreamfusion yields excessive variance and blurry results for standard CFG values, while using higher CFG values leads to over-saturation and lack of detail. On the other hand, solving eq. 8 exactly is challenging due to its high dimensionality and nonlinearity.
@@
We suggest to obtain $\kappa^{t}$ by inverting DDIM, that is, by integrating the ODE in eq. 1 with $t$ evolving backwards (from images to noise) [song2020denoising, mokady2023null, miyake2023negative]. This process approximates but is not identical to the exact solution for $\kappa^{t}$: fixed points of eq. 8 invert a single large step of DDIM, while running the ODE in reverse inverts the entire DDIM trajectory. We ablate alternative choices for $\kappa^{t}$ in section 6.2 and conclude that in practice DDIM inversion offers the best approximation quality. Additionally, to match the iterative nature of DDIM, we employ a linear annealing schedule of $t$. We refer to our modified version of SDS as Score Distillation via Inversion, or SDI.
@@
Figure 6: Comparison of intermediate variables in SDS and SDI for different timesteps $t$. Starting with a rendering of a 3D shape we demonstrate how each algorithm perturbs it ($x(t)$ variable on the top row) and how it is denoised with a single step of diffusion ($x_{0}(t)$ variable on the bottom row). The prompt used is “Pumpkin head zombie, skinny, highly detailed, photorealistic, side view.”
@@
Figure 6 shows the effect of inferring the noise via DDIM inversion instead of sampling it randomly. The special structure of the improved $\kappa^{t}$ results in more consistent single-step generations and produces intricate features at earlier times. When inverted and not sampled, the noise appears ‘in the right place’: in SDS the noise covers the whole view, including the background, whereas in ours the noise is concentrated on the meaningful part of the 3D shape. This improves geometric and temporal coherency for $x_{0}(t)$ predictions even for large $t$. The reduced variance drastically increases sharpness and level of detail. Moreover, it allows to reduce CFG value of generation $\gamma_{\text{fwd}}$ to the standard $7.5$, avoiding oversaturation. Another interesting finding is that DDIM inversion works best when the reverse integration is performed with negative CFG $\gamma_{\text{inv}}=-\gamma_{\text{fwd}}=-7.5$. The overview of our method is presented in Figure 5, the details about inversion algorithm are presented in section 6.2, and the implementation details are discussed in appendix A.
@@ 6 Experiments
We demonstrate the high-fidelity 3D shapes generated with our algorithm in fig. 2, provide more examples of $360^{\circ}$ views in section E.2 and videos in the supplementary materials.
Qualitative comparisons. Figure 7 compares 3D generation quality with the results reported in past work using a similar protocol to [wang2024prolificdreamer, katzir2023noise]. For the baselines we chose: Dreamfusion [poole2022dreamfusion] (the work we build on), Noise Free Score Distillation (NFSD) [katzir2023noise] (uses negative prompts in SDS), ProlificDreamer [wang2024prolificdreamer] (fine-tunes and trains a neural network to denoise the 3D shape), and HiFA [zhu2023hifa] (guides in both image and latent spaces, regularizes the NeRF, and supervises the geometry with mono-depth estimation). Our figures indicate that Score Distillation via Inversion (SDI) yields similar or better results compared with state-of-the-art. Section E.3 presents more extensive comparison.
Figure 7: Comparison of 3D generation with other methods using their reported results. The prompts are “A 3D model of an adorable cottage with a thatched roof” and “A car made out of sushi.” The right side provides 2D images generated with the same diffusion model via DDIM for reference.
@@ 6.2 Ablations
Figure 8: Ablation study of proposed improvements.
Choice of $\kappa^{t}(x)$. The key component of our algorithm is the choice of inferred noise $\kappa^{t}(x)$. In theory, $\kappa^{t}(x)$ should solve eq. 8, but it is impractical to do so. Hence, in fig. 9 we compare the following choices of $\kappa^{t}(x)$ and report their numerical errors:
  • Random, re-sampled: Sample $\kappa^{t}(x)$ in each new update step from $\mathcal{N}(0,I)$ ;
  • Random, fixed: Sample $\kappa^{t}(x)$ from $\mathcal{N}(0,I)$ once, and keep it fixed for each iteration;
  • Fixed point iteration: Since the optimal solution is a fixed point of eq. 8, initialize $\kappa^{t}(x)\sim\mathcal{N}(0,I)$ and run fixed point iteration [burden19852] for 10 steps (in practice, more steps did not help).
  • SGD optimization: Optimize $\kappa^{t}(x)$ via gradient decent for 10 steps, initializing with noise.
@@
The left side of fig. 9 compares the choices for 3D generation qualitatively; the right side plots error induced in eq. 8 (re-scaled to $x_{0}$ variable due to its ambiguity around 00). As can be seen, the regressed noise has a big impact on the final generations. Both fixed point iteration and optimization via gradient descent fail to improve the approximation of $\kappa^{t}$, while fixed point iteration diverges in 3D. On the other hand, DDIM inversion yields a reasonable approximation of $\kappa^{t}$ and significantly improves 3D generation quality.
Figure 9: The ablation study of different $\kappa^{t}$ choices in our algorithm. We show the obtained 3D generations on the left (Fixed Point Iteration diverges), and the numerical error in eq. 8 induced for each timestep on the right. The resampled and fixed noise strategies produce the same error.
CFG for inversion. [mokady2023null, miyake2023negative] report that DDIM inversion accumulates big numerical error for CFG $\gamma>1$. Surprisingly, we find that DDIM inversion for CFG $\gamma_{\text{fwd}}>1$ can be adequately estimated by running the inversion with negative CFG $\gamma_{\text{inv}}=-\gamma_{\text{fwd}}$. Figure 10 compares 3D inversion strategies qualitatively and quantitatively. Naïvely taking $\gamma_{\text{inv}}=\gamma_{\text{fwd}}=7.5$ yields the biggest numerical error, while other strategies perform on par. 3D generations, illustrated on the right, show that $\gamma_{\text{inv}}=\gamma_{\text{fwd}}=7.5$ introduces excessive numerical errors, causing generation to drift in a random direction. The best parameters (as we demonstrate in appendix B) for 2D inversion ($\gamma_{\text{inv}}=\gamma_{\text{fwd}}=1$) fail to converge in 3D as there is not enough guidance toward a class sample. Introducing guidance only on the forward pass with $\gamma_{\text{inv}}=1,\gamma_{\text{fwd}}=7.5$ solves the problem, and the algorithm generates the desired 3D shape, but constantly adding CFG on each step over-saturates the image. Using $\gamma_{\text{inv}}=-7.5,\gamma_{\text{fwd}}=7.5$ cancels the over-saturation and produces accurate 3D generations.

### v2

@@ 1 Introduction
Based on our analysis, we propose an alternative score distillation algorithm dubbed Score Distillation via Inversion (SDI), closing the gap to DDIM. We obtain the conditional noise required for consistency of the denoising trajectories by inverting DDIM on each step of score distillation (fig. 5). This modification yields 3D objects with high-quality textures consistent with the 2D diffusion model (fig. 1e). Moreover, in 2D, our method closely approximates DDIM while preserving the incremental generation schedule of SDS (fig. 1c).
@@
  • We propose a new method titled Score Distillation via Inversion (SDI), which replaces the problematic random noise sampling in SDS with prompt-conditioned DDIM inversion and significantly improves 3D generation, closing the quality gap to samples from the 2D model.
@@ 2 Related work
Other papers propose heuristics that improve SDS. For example, [katzir2023noise, yu2023text, tang2023stable] decompose the guidance terms and speculate about their relative importance. Empirically, visual quality can be improved by suppressing the denoising term with negative prompts [katzir2023noise] or highlighting the classification term [yu2023text]. [lin2023magic3d, tang2023make, wang2024prolificdreamer, chen2024it3d] use multi-stage optimization: they first train a volumetric representation and then extract a mesh or voxel grid to fine-tune geometry and texture. HiFA [zhu2023hifa] uses a combination of ad-hoc techniques: additional supervision in latent space, time annealing, kernel smoothing, $z$-variance regularization, and a pretrained mono-depth estimation network to improve the quality of single-stage NeRF generation.
LucidDreamer or Interval Score Matching (ISM), hypothesizes based on empirical evidence that in SDS, the high-variance random noise term, together with the big reconstruction error of a single-step prediction cause over-smoothing [liang2023luciddreamer]. Based on these observations, the authors replace the random noise term in SDS with the noise obtained by running DDIM inversion process, and introduce multi-step denoising to improve the reconstruction quality. As we will show in section 4, the update rule of ISM is closely related to ours and can be seen as a special case of our formulation. Moreover, our analysis reveals that the added noise should be inferred conditionally on the text prompt $y$, which further leads to quality improvement.
In this work, rather than augmenting the SDS pipeline or relying on heuristics, we derive Score Distillation through the denoising process of DDIM and propose a simple modification of SDS that significantly improves 3D generation.
@@ 3 Background
Denote the parameters of a differentiable 3D shape representation by $\psi\in\mathbb{R}^{d}$ and differentiable rendering by a function $g(\psi,c):\mathbb{R}^{d}\times C\rightarrow\mathbb{R}^{N\times N}$ that returns an image given camera parameters $c\in C$. Intuitively, in each iteration, SDS samples $c$, renders the corresponding (image) view $g(\psi,c)$, perturbs it with $\epsilon\sim\mathcal{N}(0,I)$ to level $t\sim[0,1]$, and denoises it with $\epsilon_{\theta}^{t}$; the difference between the true and predicted noise is propagated to the parameters of the 3D shape. More formally, after sampling the camera view $c$ and randomly drawing a time $t$, SDS renders the volume and adds Gaussian noise $\epsilon$ to obtain a noisy image $x(t)=\sqrt{\alpha(t)}g(\psi,c)+\sqrt{1-\alpha(t)}\epsilon.$ Then, SDS improves the generated volume by using a gradient(-like) direction to update its parameters $\psi$:
@@ 4 Linking SDS to DDIM
In words, $x_{0}(t)$ is the noisy image at time $t$ denoised with a single step of noise prediction. Empirically, the evolution of $x_{0}(t)$ is similar to the evolution of $g(\psi,c)$—from blurry to sharp. In the left side of fig. 4 we compare these processes. This similarity motivates us to rewrite eq. 1 in terms of $x_{0}(t)$, and to understand SDS as applying similar updates to the renderings of its 3D representation.
@@
Define $\kappa^{t}_{y}\bigl{(}x_{0}(t)\bigr{)}=\epsilon$ as a solution of this equation given $x_{0}(t)$. Then, we can write:
$\epsilon_{\theta}^{t}\bigl{(}x(t),y\bigr{)}=\kappa^{t}_{y}\bigl{(}x_{0}(t)\bigr{)}\textrm{\quad and\quad}x(t-\tau)=\sqrt{\alpha(t-\tau)}x_{0}(t)+\sqrt{1-\alpha(t-\tau)}\kappa^{t}_{y}\bigl{(}x_{0}(t)\bigr{)}\text{.}$ (9)
@@
$x_{0}(t\!-\!\tau)=x_{0}(t)\!-\!\sigma(t\!-\!\tau)\bigl{[}\underbrace{\epsilon_{\theta}^{t-\tau}\bigl{(}\overbrace{\sqrt{\alpha(t\!-\!\tau)}x_{0}(t)\!+\!\sqrt{1\!-\!\alpha(t\!-\!\tau)}\kappa^{t}_{y}\bigl{(}x_{0}(t)\bigr{)}}^{\text{$x_{0}$ noised with $\kappa^{t}_{y}$ to time $t-\tau$}},y\bigr{)}}_{\text{predicted noise}}\!-\!\underbrace{\kappa^{t}_{y}\bigl{(}x_{0}(t)\bigr{)}}_{\text{noise sample $\kappa^{t}_{y}$}}\!\!\!\bigr{]}\!\text{.}$ (10)
We can already see that the structure of eq. 10 is very similar to the SDS update rule in eq. 4. Note that the update direction in eq. 10 is the same as in the SDS update rule in eq. 4, where $\kappa^{t}_{y}$ plays the role of the random noise sample $\epsilon$. We could use it as a guidance for the 3D generative process in SDS by replacing $\epsilon$ in eq. 4 with $\kappa^{t}_{y}(x_{0}(t)).$ In practice, however, it is hard to solve eq. 8, as $\epsilon_{\theta}^{t}$ is high-dimensional and nonlinear. In an unconstrained 2D generation, $\kappa^{t}_{y}$ can be cached from a previous denoising step, matching the update step to DDIM exactly as in fig. 1c. In 3D, however, this is impossible due to the simultaneous optimization of multiple views and projections to the space of viable 3D shapes. Below we show that a naïve approximation replacing $\kappa^{t}_{y}$ with a Gaussian yields SDS, and we will propose alternatives that are more faithful to the derivation above.
SDS as a special case. From eq. 10, to get a cleaner image, we need to bring the current image to time $t$ with noise sample $\kappa^{t}_{y}$, denoise the obtained image, and then subtract the difference between added and predicted noise from the initial image. A coarse approximation of $\kappa^{t}_{y}$ uses i.i.d. random noise $\kappa_{\textrm{SDS}}(x_{0}(t))\sim\mathcal{N}(0,I)$, matching the forward process by which diffusion adds noise. This choice of $\kappa_{\textrm{SDS}}$ precisely matches the update rule eq. 10 to the SDS guidance in eq. 4.
ISM as a special case. The main update formula in Interval Score Matching (ISM) [liang2023luciddreamer] is a particular case of eq. 10, where $\kappa^{t}_{y}$ is obtained via DDIM inversion without conditioning on the text prompt $y$. As demonstrated in section 5, DDIM inversion approximates a solution of eq. 10, which explains the improved performance of ISM. However, our analysis suggests that an even better-performing $\kappa^{t}_{y}$ should incorporate the text prompt $y$, which further improves the results and allows to avoid over-saturation.
@@ 5 Score Distillation via Inversion (SDI)
As we have shown, SDS follows the velocity field of reparametrized DDIM in eq. 10, when $\kappa_{\textrm{SDS}}(x_{0}(t))$ is randomly sampled in each step. Our derivation, however, suggests that $\kappa_{\textrm{SDS}}(x_{0}(t))$ could be improved by bringing it closer to a solution of the fixed-point equation in eq. 8. Indeed, randomly sampling $\kappa_{\textrm{SDS}}$ as in Dreamfusion yields excessive variance and blurry results for standard CFG values, while using higher CFG values leads to over-saturation and lack of detail. On the other hand, solving eq. 8 exactly is challenging due to its high dimensionality and nonlinearity.
@@
Like ISM [liang2023luciddreamer], we suggest to obtain $\kappa^{t}_{y}$ by inverting DDIM, that is, by integrating the ODE in eq. 1 with $t$ evolving backwards (from images to noise) as in [song2020denoising, mokady2023null, miyake2023negative]. As we can see in eq. 8, $\kappa^{t}_{y}$ should be a function of the text prompt $y$, leading us to perform DDIM inversion conditionally on $y$, unlike ISM. This process approximates but is not identical to the exact solution for $\kappa^{t}_{y}$: fixed points of eq. 8 invert a single large step of DDIM, while running the ODE in reverse inverts the entire DDIM trajectory. We ablate alternative choices for $\kappa^{t}_{y}$ in section 6.2 and conclude that in practice DDIM inversion offers the best approximation quality. Additionally, to match the iterative nature of DDIM, we employ a linear annealing schedule of $t$. We refer to our modified version of SDS as Score Distillation via Inversion, or SDI.
@@
Figure 6: Comparison of intermediate variables in SDS and SDI (ours) for different timesteps $t$. Starting with a rendering of a 3D shape we demonstrate how each algorithm perturbs it ($x(t)$ variable on the top row) and how it is denoised with a single step of diffusion ($x_{0}(t)$ variable on the bottom row). The prompt used is “Pumpkin head zombie, skinny, highly detailed, photorealistic, side view.”
@@
Figure 6 shows the effect of inferring the noise via DDIM inversion instead of sampling it randomly. The special structure of the improved $\kappa^{t}_{y}$ results in more consistent single-step generations and produces intricate features at earlier times. When inverted and not sampled, the noise appears ‘in the right place’: in SDS the noise covers the whole view, including the background, whereas in ours the noise is concentrated on the meaningful part of the 3D shape. This improves geometric and temporal coherency for $x_{0}(t)$ predictions even for large $t$. The reduced variance drastically increases sharpness and level of detail. Moreover, it allows to reduce CFG value of generation $\gamma_{\text{fwd}}$ to the standard $7.5$, avoiding over-saturation. Another interesting finding is that DDIM inversion works best when the reverse integration is performed with negative CFG $\gamma_{\text{inv}}=-\gamma_{\text{fwd}}=-7.5$. The overview of our method is presented in fig. 5, the details about inversion algorithm are presented in section 6.2, and the implementation details are discussed in appendix A.
@@ 6 Experiments
We demonstrate the high-fidelity 3D shapes generated with our algorithm in fig. 2 and provide more examples of $360^{\circ}$ views in section F.2.
Qualitative comparisons. Figure 7 compares 3D generation quality with the results reported in past work using a similar protocol to [wang2024prolificdreamer, katzir2023noise]. For the baselines we chose: Dreamfusion [poole2022dreamfusion] (the work we build on), Noise Free Score Distillation (NFSD) [katzir2023noise] (uses negative prompts in SDS), ProlificDreamer [wang2024prolificdreamer] (fine-tunes and trains a neural network to denoise the 3D shape), Interval Score Matching (ISM) [liang2023luciddreamer] (obtain the noise sample by inverting DDIM and perform multi-step denoising), and HiFA [zhu2023hifa] (guides in both image and latent spaces, regularizes the NeRF, and supervises the geometry with mono-depth estimation). Our figures indicate that Score Distillation via Inversion (SDI) yields similar or better results compared with state-of-the-art. Section F.3 presents more extensive comparison.
Figure 7: Comparison of 3D generation with other methods using their reported results. The prompts are “An ice cream sundae” and “A 3D model of an adorable cottage with a thatched roof”.
@@ 6.2 Ablations
Choice of $\kappa^{t}_{y}(x)$. The key component of our algorithm is the choice of inferred noise $\kappa^{t}_{y}(x)$. In theory, $\kappa^{t}_{y}(x)$ should solve eq. 8, but it is impractical to do so. Hence, in fig. 9 we compare the following choices of $\kappa^{t}_{y}(x)$ and report their numerical errors:
  • Random, re-sampled: Sample $\kappa^{t}_{y}(x)$ in each new update step from $\mathcal{N}(0,I)$ ;
  • Random, fixed: Sample $\kappa^{t}_{y}(x)$ from $\mathcal{N}(0,I)$ once, and keep it fixed for each iteration;
  • Fixed point iteration: Since the optimal solution is a fixed point of eq. 8, initialize $\kappa^{t}_{y}(x)\sim\mathcal{N}(0,I)$ and run fixed point iteration [burden19852] for 10 steps (in practice, more steps did not help).
  • SGD optimization: Optimize $\kappa^{t}_{y}(x)$ via gradient decent for 10 steps, initializing with noise.
@@
The left side of fig. 9 compares the choices for 3D generation qualitatively; the right side plots error induced in eq. 8 (re-scaled to $x_{0}$ variable due to its ambiguity around 00). As can be seen, the regressed noise has a big impact on the final generations. Both fixed point iteration and optimization via gradient descent fail to improve the approximation of $\kappa^{t}_{y}$, while fixed point iteration diverges in 3D. On the other hand, DDIM inversion yields a reasonable approximation of $\kappa^{t}_{y}$ and significantly improves 3D generation quality. We provide visual comparison of the obtained noisy images for each baseline in appendix B.
Figure 9: The ablation study of different $\kappa^{t}_{y}$ choices in our algorithm. We show the obtained 3D generations on the left (Fixed Point Iteration diverges), and the numerical error in eq. 8 induced for each timestep on the right. The resampled and fixed noise strategies produce the same error.
CFG for inversion. [mokady2023null, miyake2023negative] report that DDIM inversion accumulates big numerical error for CFG $\gamma>0$. Surprisingly, we find that DDIM inversion for CFG $\gamma_{\text{fwd}}>0$ can be adequately estimated by running the inversion with negative CFG $\gamma_{\text{inv}}=-\gamma_{\text{fwd}}$. Figure 10 compares 3D inversion strategies qualitatively and quantitatively. Naïvely taking $\gamma_{\text{inv}}=\gamma_{\text{fwd}}=7.5$ yields the biggest numerical error, while other strategies perform on par. 3D generations, illustrated on the right, show that $\gamma_{\text{inv}}=\gamma_{\text{fwd}}=7.5$ introduces excessive numerical errors, causing generation to drift in a random direction. The best parameters (as we demonstrate in appendix C) for 2D inversion ($\gamma_{\text{inv}}=\gamma_{\text{fwd}}=0$) fail to converge in 3D as there is not enough guidance toward a class sample. Introducing guidance only on the forward pass with $\gamma_{\text{inv}}=0,\gamma_{\text{fwd}}=7.5$ solves the problem, and the algorithm generates the desired 3D shape, but constantly adding CFG on each step over-saturates the image. Note, in that configuration the inversion is performed unconditionally from the text prompt, matching ISM [liang2023luciddreamer]. As we can see, prompt conditioning is an important component in eq. 8, and using $\gamma_{\text{inv}}=-7.5,\gamma_{\text{fwd}}=7.5$ cancels the over-saturation and produces accurate 3D generations.